# Quantum Walks and Exact RG in de Sitter Space

### v1

@@ 1 Introduction
Recently, it has been shown that the random walk description of light scalars can be recovered in QFT on a fixed dS background using RG techniques [8, 9, 10, 11, 12, 13, 14, 15, 16]. In the process, many of these long standing issues are resolved at both a technical and conceptual level. From the RG point of view, the theory evolves away from the trivial fixed point to a non-trivial self-similar description that is equivalent to a random walk. At the same time, this description must have a limited regime of validity; a classical random walk generates entropy which cannot be the case in a closed quantum system. The stochastic formalism is therefore limited to subsystems within de Sitter space [17, 18, 19], although that aspect is not always transparent in the derivation. A complete understanding of light fields in de Sitter should furnish an explanation for the origin and breakdown of this description.
@@ 3 Exact RG and the Density Matrix
We have intuition from RG flow and EFT that suggests that the evolution of the reduced density matrix can be understood as the exact RG for the density matrix $\rho[\phi_{+},\phi_{-}]$ after integrating our a shell of modes with $k\in[\Lambda,\Lambda+\delta\Lambda]$ where $\Lambda>aH$. For purely long wavelength observables, we are going to trace over all the short modes anyway and $\rho_{R}$ is independent of $\Lambda\gg aH$, and plays the same role as the Euclidean partition function of an exact RG in flat space [38]. However, because of the expansion of the Universe, modes that are in $\rho_{R}(t)$ would not have been in $\rho_{R}(t^{\prime})$ for $t^{\prime}\ll t$. Therefore, understanding dynamical RG (RG plus time evolution) introduces subtleties beyond the usual exact RG. Nevertheless, we will find that the by following analogous exact RG treatment for the density matrix in flat space [39, 31], we can derive the evolution of the density matrix in dS under time evolution.
@@ 3.2 Integrating Out
Using the obvservation that
@@
However, since the local action can only in involve operators in one of the $\gamma_{+}=t_{+}\in(-\infty(1-i\epsilon),t_{0}]$ or $\gamma_{-}=t_{-}\in[t_{0},-\infty(1+i\epsilon)]$ region, we may still have terms mixing $t_{+}$ and $t_{-}$ that do not arise in the bulk action, see Figure 1 for illustration. Therefore, defining
@@
While the presence of the additional boundary terms will be important for cosmological applications, such as stochastic inflation, it is important to note that they are needed in flat space when matching correlators [31]. The fact they do not appear in matching S-matrix elements is a consequence of the fact that they can represent field redefinitions between the UV and IR descriptions.
@@
where $\mathdollar_{\rm int}$ are higher order local terms defined in Equation (3.34). We can further simplify the quadratic terms in this equation equation by recognizing that
@@ 4.1 Leading Order
There are, of course, easier ways to arrive at this result. What is meaningful here is that the determination of this PDE from the exact RG as a functional over all the Fourier models is more nuanced than one might have expected from just comparing Equations (3.38) and (3.39).
@@ 4.2 Next-to-Leading Order
For $V(\bar{\varphi})=\frac{\lambda{\bar{\varphi}}^{4}}{4!}$, if we interpret $3Hq=-V_{\rm eff}^{\prime}$, the we can see that
@@
$\frac{d}{dt}W[\bar{\varphi},q]$ $=3HW+\frac{H^{3}}{8\pi^{2}}\frac{\partial^{2}W}{\partial{\bar{\varphi}}^{2}}+a^{3}V_{\mathrm{eff}}^{\prime\prime}(\varphi)\frac{H^{2}}{12\pi^{2}}\frac{\partial^{2}W}{\partial Q\partial{\bar{\varphi}}}+a^{6}\frac{H}{72\pi^{2}}(V_{\mathrm{eff}}^{\prime\prime})^{2}\frac{\partial^{2}W}{\partial Q^{2}}$
@@ 4.3 NNLO and Beyond
An essential difference between the approach in this paper and [59] is that we are working in a Wilsonsian EFT, rather than a continuum EFT [62]. In the Wilsonian description, we integrated out the short modes explicitly, giving rise to higher dimension interactions, including higher derivative terms, in $S_{\rm eff}$. These terms do not arise in a continuum theory, as they would vanish in dimensional regularization. Nevertheless, the corrections to these terms in the Wilsonian description are needed to reproduce higher order contributions to the RG flow [38]. In this regard, the challenge of calculating NNLO corrections in our exact RG scheme is not unique to de Sitter, but a common aspect of Wilsonian versus continuum descriptions.
@@ 5 Renomalization in dS
In flat slicing of dS, the Weyl transformation that takes flat space correlators to cosmological correlators involves scaling by the scale factor $a(\tau)$. For a CFT, this requires that any anomalous dimensions are associated with time-dependent logs. The required matching between dS and flat space therefore strongly suggests that the RG resummation involves time-dependent logs in dS. In this sense, it is dynamical RG that should be on the same footing as conventional RG in flat space [^8] For principle series fields, some anomalous dimensions can be calculated by resumming bubble diagrams, more like a decay width [71]. These examples are compatible with the RG description, as shown in [72]. Yet, the conformal example [49] demonstrates that not all anomalous dimensions can be calculated by summing geometric series..
@@ B.1 Identity useful for homogenization
This is, of course, only true at the Gaussian level. Now, we can evaluate terms of the form
@@ Appendix C Eliminating Momentum with Moments
$\frac{d}{dt}W[\bar{\varphi},Q]$ $=\frac{H^{3}}{8\pi^{2}}\frac{\partial^{2}W}{\partial\varphi^{2}}+a^{3}V_{\mathrm{eff}}^{\prime\prime}(\varphi)\frac{H^{2}}{12\pi^{2}}\frac{\partial^{2}W}{\partial Q\partial\varphi}+a^{6}\frac{H}{72\pi^{2}}(V_{\mathrm{eff}}^{\prime\prime})^{2}\frac{\partial^{2}W}{\partial Q^{2}}$
$-V_{\mathrm{eff}}^{\prime\prime}\frac{\partial}{\partial Q}\left(\frac{Q}{3H}W\right)-\frac{Q}{a^{3}}\frac{\partial W}{\partial\varphi}+\frac{1}{3H}\frac{\partial}{\partial\varphi}(V_{\mathrm{eff}}^{\prime}W)\ ,$ (C.2)

### v2

@@ 1 Introduction
Recently, it has been shown that the random walk description of light scalars can be recovered in QFT on a fixed dS background using RG techniques [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In the process, many of these long standing issues are resolved at both a technical and conceptual level. From the RG point of view, the theory evolves away from the trivial fixed point to a non-trivial self-similar description that is equivalent to a random walk. At the same time, this description must have a limited regime of validity; a classical random walk generates entropy which cannot be the case in a closed quantum system. The stochastic formalism is therefore limited to subsystems within de Sitter space [18, 19, 20], although that aspect is not always transparent in the derivation. A complete understanding of light fields in de Sitter should furnish an explanation for the origin and breakdown of this description.
@@ 3 Exact RG and the Density Matrix
We have intuition from RG flow and EFT that suggests that the evolution of the reduced density matrix can be understood as the exact RG for the density matrix $\rho[\phi_{+},\phi_{-}]$ after integrating our a shell of modes with $k\in[\Lambda,\Lambda+\delta\Lambda]$ where $\Lambda>aH$. For purely long wavelength observables, we are going to trace over all the short modes anyway and $\rho_{R}$ is independent of $\Lambda\gg aH$, and plays the same role as the Euclidean partition function of an exact RG in flat space [39]. However, because of the expansion of the Universe, modes that are in $\rho_{R}(t)$ would not have been in $\rho_{R}(t^{\prime})$ for $t^{\prime}\ll t$. Therefore, understanding dynamical RG (RG plus time evolution) introduces subtleties beyond the usual exact RG. Nevertheless, we will find that the by following analogous exact RG treatment for the density matrix in flat space [40, 32] (see also [52, 53]), we can derive the evolution of the density matrix in dS under time evolution. Our approach is related to previous work deriving the Stochastic formalism using functional/exact RG [54, 55, 56, 57].
@@ 3.2 Integrating Out
Using the observation that
@@
However, since the local action can only involve operators in one of the $\gamma_{+}=t_{+}\in(-\infty(1-i\epsilon),t_{0}]$ or $\gamma_{-}=t_{-}\in[t_{0},-\infty(1+i\epsilon)]$ regions, we may still have terms mixing $t_{+}$ and $t_{-}$ that do not arise in the bulk action, see Figure 1 for illustration. Therefore, defining
@@
While the presence of the additional boundary terms will be important for cosmological applications, such as stochastic inflation, it is important to note that they are needed in flat space when matching correlators [32]. The fact that they do not appear in matching S-matrix elements is a consequence of the fact that they can represent field redefinitions between the UV and IR descriptions.
@@
where $\mathdollar_{\rm int}$ are higher order local terms defined in Equation (3.34). We can further simplify the quadratic terms in this equation by recognizing that
@@ 4.1 Leading Order
There are, of course, easier ways to arrive at this result. What is meaningful here is that the determination of this PDE from the exact RG as a functional over all the Fourier modes is more nuanced than one might have expected from just comparing Equations (3.38) and (3.39).
@@ 4.2 Next-to-Leading Order
For $V(\bar{\varphi})=\frac{\lambda{\bar{\varphi}}^{4}}{4!}$, if we interpret $3Hq=-V_{\rm eff}^{\prime}$, then we can see that
@@
$\frac{d}{dt}W[\bar{\varphi},q]$ $=3HW+\frac{H^{3}}{8\pi^{2}}\frac{\partial^{2}W}{\partial{\bar{\varphi}}^{2}}+a^{3}V_{\mathrm{eff}}^{\prime\prime}\frac{H^{2}}{12\pi^{2}}\frac{\partial^{2}W}{\partial Q\partial{\bar{\varphi}}}+a^{6}\frac{H}{72\pi^{2}}(V_{\mathrm{eff}}^{\prime\prime})^{2}\frac{\partial^{2}W}{\partial Q^{2}}$
@@ 4.3 NNLO and Beyond
An essential difference between the approach in this paper and [66] is that we are working in a Wilsonian EFT, rather than a continuum EFT [69]. In the Wilsonian description, we integrated out the short modes explicitly, giving rise to higher dimension interactions, including higher derivative terms, in $S_{\rm eff}$. These terms do not arise in a continuum theory, as they would vanish in dimensional regularization. Nevertheless, the corrections to these terms in the Wilsonian description are needed to reproduce higher order contributions to the RG flow [39]. In this regard, the challenge of calculating NNLO corrections in our exact RG scheme is not unique to de Sitter, but a common aspect of Wilsonian versus continuum descriptions.
@@ 5 Renomalization in dS
In flat slicing of dS, the Weyl transformation that takes flat space correlators to cosmological correlators involves scaling by the scale factor $a(\tau)$. For a CFT, this requires that any anomalous dimensions are associated with time-dependent logs. The required matching between dS and flat space therefore strongly suggests that the RG resummation involves time-dependent logs in dS. In this sense, it is dynamical RG that should be on the same footing as conventional RG in flat space [^8] For principle series fields, some anomalous dimensions can be calculated by resumming bubble diagrams, more like a decay width [78, 79]. These examples are compatible with the RG description, as shown in [80, 81]. Yet, the conformal example [50] demonstrates that not all anomalous dimensions can be calculated by summing geometric series..
@@ B.1 Identity useful for homogenization
This is, of course, only true at the Gaussian level, which is sufficient for our purposes right now. Now, we can evaluate terms of the form
@@ Appendix C Eliminating Momentum with Moments
$\frac{d}{dt}W[\bar{\varphi},Q]$ $=\frac{H^{3}}{8\pi^{2}}\frac{\partial^{2}W}{\partial\varphi^{2}}+a^{3}V_{\mathrm{eff}}^{\prime\prime}({\bar{\varphi}})\frac{H^{2}}{12\pi^{2}}\frac{\partial^{2}W}{\partial Q\partial{\bar{\varphi}}}+a^{6}\frac{H}{72\pi^{2}}(V_{\mathrm{eff}}^{\prime\prime}({\bar{\varphi}}))^{2}\frac{\partial^{2}W}{\partial Q^{2}}$
$-V_{\mathrm{eff}}^{\prime\prime}({\bar{\varphi}})\frac{\partial}{\partial Q}\left(\frac{Q}{3H}W\right)-\frac{Q}{a^{3}}\frac{\partial W}{\partial{\bar{\varphi}}}+\frac{1}{3H}\frac{\partial}{\partial{\bar{\varphi}}}(V_{\mathrm{eff}}^{\prime}({\bar{\varphi}})W)\ ,$ (C.2)